OpenStudy (anonymous):

Calculate the distance from the point (3,5) to the line y = x + 4.

OpenStudy (anonymous):

choose a point on y = x+ 4 (0,4) now use distance formula for two points of (0,4) and (3,5)

OpenStudy (anonymous):

 The shortest distance will be along a line that is perpendicular to the given line. That line will have the opposite, reciprocal slope, so it would be y = -x +b. To find its b value, plug in the point's coordinates:\[5 = -3+b\]\[b =8\] So the perpendicular line that passes through the point is y = -x+8. To find the location on the original line that is closest to the given point, find the intersection between the two lines:\[x+4 = -x+8\]\[2x = 4\]\[x =2\] Then plug in 2 for x in either equation to find the y value:\[y = x+ 4 = 2+4 = 6\] So the closest point on the original line to the given point is (2, 6). Now use the distance formula to find the distance between the two points\[d = \sqrt{(\Delta x)^{2} + (\Delta y)^{2}} = \sqrt{(3-2)^{2} + (5-6)^{2}} = \sqrt{2}\]